# StyleSinger 2: Zero-Shot Singing Voice Synthesis with Style Transfer and Multi-Level Style Control

### v1

@@ 4.3 Ablation Study
As depicted in Table 5, we conduct ablation studies to showcase the efficacy of various designs within StyleSinger 2. We use CMOS-Q to test the variation in synthesis quality, and CMOS-S to measure the changes in singer similarity. 1) Using VQ instead of CVQ in the clustering style encoder resulted in decreased synthesis quality and singer similarity, indicating the importance of CVQ for stable and high-quality style extraction. 2) Eliminating the style adaptive decoder and using an 8-step diffusion decoder (Huang et al., 2022b) led to declines in both synthesis quality and singer similarity, underscoring the role of our method in enhancing style diversity in singing voices. 3) Predicting only styles in the S&D-LM while using a simple duration predictor (Ren et al., 2020) for phoneme duration also resulted in decreased synthesis quality and singer similarity. This demonstrates the mutual benefits of our model in predicting both phoneme duration and style information. Please refer to Appendix E.1 for more results.
@@ 5 Conclusion
In this paper, we introduce StyleSinger 2, the first zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. StyleSinger 2 transfers and controls styles (like singing methods, emotion, rhythm, technique, and pronunciation) from audio and text prompts to synthesize high-quality singing voices. The performance of our model is primarily enhanced through three key components: 1) the clustering style encoder that stably condenses style information into a compact latent space using a CVQ model, thus facilitating subsequent predictions; 2) the Style and Duration Language Model (S&D-LM), which predicts style information and phoneme duration simultaneously, which benefits both; and 3) the style adaptive decoder that employs a novel mel-style adaptive normalization method to generate enhanced details in singing voices. Experimental results demonstrate that StyleSinger 2 surpasses baseline models in synthesis quality, singer similarity, and style controllability across zero-shot style transfer, multi-level style control, cross-lingual style transfer, and STS style transfer.
@@ 7 Ethics Statement
StyleSinger 2, with its capability to transfer and control diverse styles of singing voices, could potentially be misused for dubbing in entertainment videos, raising concerns about the infringement of singers’ copyrights. Additionally, its ability to transfer cross-lingual speech and singing styles poses risks of unfair competition and potential unemployment for professionals in related singing occupations. To mitigate these risks, we will implement stringent restrictions on the use of our model to prevent unauthorized and unethical applications. We will also explore methods such as vocal watermarking to protect individual privacy.
@@ Acknowledgements
This work was supported in part by the National Key R&D Program of China under Grant No.2022ZD0162000, National Natural Science Foundation of China under Grant No.62222211, Grant No.61836002 and Grant No.62072397, and Yiwise.
@@ A.1 Architecture Details
We list the architecture and hyperparameters of our StyleSinger 2 in Table 6.
@@
Table 6: Hyper-parameters of StyleSinger 2 modules.
@@ Appendix B Details of Dataset
For phonetic content, Chinese phonemes were extracted using pypinyin  [^1] https://github.com/mozillazg/python-pinyin, English phonemes followed the ARPA standard  [^2] https://en.wikipedia.org/wiki/ARPABET. We selected these standards because Chinese uses pinyin for pronunciation and ARPA includes English stress patterns, making them the most suitable phoneme standards for each language. We then add all phonemes in a unified phoneme set. This strategy allows our model to embed phonemes for all languages during training for all tasks. Subsequently, we randomly chose 40 singers as the unseen test set to evaluate StyleSinger 2 in the zero-shot scenario for all tasks. Notably, our dataset partitioning carefully ensures that both training and test sets for all tasks include cross-lingual singing and speech data.
@@ E.2 Multi-Level Style Control
Method StyleSinger StyleSinger 2 (ours)
@@
method StyleSinger StyleSinger 2 (ours)
@@
Method StyleSinger StyleSinger 2 (ours)
@@
As shown, our StyleSinger 2 outperforms baseline models in style control tasks for any type of controllable style based on objective metrics.

### v2

@@ 4.3 Ablation Study
As depicted in Table 5, we conduct ablation studies to showcase the efficacy of various designs within TCSinger. We use CMOS-Q to test the variation in synthesis quality, and CMOS-S to measure the changes in singer similarity. 1) Using VQ instead of CVQ in the clustering style encoder resulted in decreased synthesis quality and singer similarity, indicating the importance of CVQ for stable and high-quality style extraction. 2) Eliminating the style adaptive decoder and using an 8-step diffusion decoder (Huang et al., 2022b) led to declines in both synthesis quality and singer similarity, underscoring the role of our method in enhancing style diversity in singing voices. 3) Predicting only styles in the S&D-LM while using a simple duration predictor (Ren et al., 2020) for phoneme duration also resulted in decreased synthesis quality and singer similarity. This demonstrates the mutual benefits of our model in predicting both phoneme duration and style information. Please refer to Appendix E.1 for more results.
@@ 5 Conclusion
In this paper, we introduce TCSinger, the first zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. TCSinger transfers and controls styles (like singing methods, emotion, rhythm, technique, and pronunciation) from audio and text prompts to synthesize high-quality singing voices. The performance of our model is primarily enhanced through three key components: 1) the clustering style encoder that stably condenses style information into a compact latent space using a CVQ model, thus facilitating subsequent predictions; 2) the Style and Duration Language Model (S&D-LM), which predicts style information and phoneme duration simultaneously, which benefits both; and 3) the style adaptive decoder that employs a novel mel-style adaptive normalization method to generate enhanced details in singing voices. Experimental results demonstrate that TCSinger surpasses baseline models in synthesis quality, singer similarity, and style controllability across zero-shot style transfer, multi-level style control, cross-lingual style transfer, and STS style transfer.
@@ 7 Ethics Statement
TCSinger, with its capability to transfer and control diverse styles of singing voices, could potentially be misused for dubbing in entertainment videos, raising concerns about the infringement of singers’ copyrights. Additionally, its ability to transfer cross-lingual speech and singing styles poses risks of unfair competition and potential unemployment for professionals in related singing occupations. To mitigate these risks, we will implement stringent restrictions on the use of our model to prevent unauthorized and unethical applications. We will also explore methods such as vocal watermarking to protect individual privacy.
@@ Acknowledgements
This work was supported by National Key R&D Program of China (2022ZD0162000).
@@ A.1 Architecture Details
We list the architecture and hyperparameters of our TCSinger in Table 6.
@@
Table 6: Hyper-parameters of TCSinger modules.
@@ Appendix B Details of Dataset
For phonetic content, Chinese phonemes were extracted using pypinyin  [^1] https://github.com/mozillazg/python-pinyin, English phonemes followed the ARPA standard  [^2] https://en.wikipedia.org/wiki/ARPABET. We selected these standards because Chinese uses pinyin for pronunciation and ARPA includes English stress patterns, making them the most suitable phoneme standards for each language. We then add all phonemes in a unified phoneme set. This strategy allows our model to embed phonemes for all languages during training for all tasks. Subsequently, we randomly chose 40 singers as the unseen test set to evaluate TCSinger in the zero-shot scenario for all tasks. Notably, our dataset partitioning carefully ensures that both training and test sets for all tasks include cross-lingual singing and speech data.
@@ E.2 Multi-Level Style Control
Method StyleSinger TCSinger (ours)
@@
method StyleSinger TCSinger (ours)
@@
Method StyleSinger TCSinger (ours)
@@
As shown, our TCSinger outperforms baseline models in style control tasks for any type of controllable style based on objective metrics.